OpenStudy (anonymous):

solve the system: y=-x^2+5x+1 and y=2x+3. please tell me how you got your answer

OpenStudy (anonymous):

you can turn it into a composite function and solve

OpenStudy (anonymous):

nah ill use the quadratic formula

OpenStudy (anonymous):

that actually will not solve it for you

OpenStudy (anonymous):

do this instead 2x +3 = x^2 +5x+1

OpenStudy (anonymous):

know what i mean?

OpenStudy (anonymous):

Use the quadratic form to get your two x's

OpenStudy (anonymous):

y = -X^2 + 5X +1 y = 2X + 3 2X + 3 = -X^2 + 5X + 1 X^2 - 3 X + 2 = 0 ( X- 1 ) ( X - 2 ) = 0 Answer X= 1 Y= 5 , or X= 2 Y= 7

OpenStudy (anonymous):

ah forgot about the negative in front of x^2

OpenStudy (mertsj):

Why did you post this question again?